OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither u=<6,-2>; v=<2,6> Please explain this concept please?

OpenStudy (kirbykirby):

To check of they are parallel: They must be scalar multiples of each other, i.e. If \[ \vec{u}=<a,b>\] then \[ \vec{v}=k\vec{u}=<ka,kb>\], for some \(k\) being a real number. To check if they are orthogonal: You can use the dot product: \[\vec{u}\cdot \vec{v}=|\vec{u}||\vec{v}|\cos\theta \], and you know they are perpendicular if the angle is 90°. So \(\cos 90°=0\), so you should expect the dot product to be 0 if they are orthogonal. (Recall, you can also calculate the dot product the following way: If \(\vec{u}=<x_1,y_1>\) and \(\vec{v}=<x_2,y_2>\), then \[ \vec{u}\cdot \vec{v}=x_1x_2+y_1y_2\]

OpenStudy (anonymous):

Thank you, but I don't understand that explanation. Can we walk thru the problem bit by bit please?

OpenStudy (kirbykirby):

So the parallel one: Say we have 2 vectors like u = <2, 4> and v = <6, 12> These are parallel since you can multiply vector u by the scalar 3 to get v 3u = <3(2), 3(4)> = <6, 12> = v So if you can divide both components and they give the same constant, they are parallel (i.e. I can divide 6/2 = 3 and 12/4 = 3... you get 3 in both cases when you divide each component respectively). Does this one make sense?

OpenStudy (anonymous):

Somewhat So using that concept, my original problem wouldn't be parallel since it can't become <2,6>? Right?

OpenStudy (kirbykirby):

yup exactly.

OpenStudy (kirbykirby):

For the orthogonal one: 2 vectors are orthogonal if they form a right angle:  So we can use the dot product to help us. One way to write the dot product is: \(\vec{u}\cdot\vec{v}=|\vec{u}||\vec{v}|\cos \theta\) So if the vectors form a right angle, that means the angle is 90° So by the dot product formula, you get: \(\vec{u}\cdot\vec{v}=|\vec{u}||\vec{v}|\cos 90° \), but \(\cos 90°=0\), so: \(\vec{u}\cdot\vec{v}=|\vec{u}||\vec{v}|(0)\\ \vec{u}\cdot\vec{v}=0\) So the dot product will be 0 if the vectors are orthogonal. But there is another equivalent formula for the dot product: If \(\vec{u}=<x_1, y_1>\) and \(\vec{v}=<x_2,y_2>\), then \[ \vec{u}\cdot\vec{v}=x_1x_2+y_1y_2\] so with your vectors, you should calculate: \(\vec{u}\cdot\vec{v}=(6)(2)+(-2)(6) \) If that gives 0, then the vectors are orthogonal

OpenStudy (anonymous):

OOOoooooooooOoOooOoo I understand the 2nd formula so much better! So it is Orthogonal! :) Correct?

OpenStudy (kirbykirby):

yes :)

OpenStudy (anonymous):

Thank you! I submitted it in and got it right :) Though on the entire exam I got a 6/10, I hope I did well enough on the essay question to boost me... :) Again, thank you <3

OpenStudy (kirbykirby):

Hopefully! :) and your welcome!